OpenStudy (anonymous):

A real estate office manages 50 apartments in a downtown building. When the rent is $900 per month, all the units are occupied. For every $25 increase in rent, one unit becomes vacant. On average, all units require $75 in maintenance and repairs each month. How much rent should the real estate office charge to maximize profits? P(x) = R(x) - C(x) This question I am stuck. I don't know how I would formulate the equations for revenue and cost. :|

OpenStudy (amistre64):

50(75) are the costs....$3750

OpenStudy (amistre64):

or is that only 75 per unit rented?

OpenStudy (amistre64):

the rate of change is linear.... 900,50 925,49..... for every 25 over we drop by 1 50=900/25+b 5=90/25+b b = 5 - 90/25 = 7/5 rooms = cost/25 +(7/5) ??

OpenStudy (amistre64):

900,50; and 925,49 50 + 900/25 = b lol i forgot to negate the slope

OpenStudy (amistre64):

50+36 = 86 y = -x/25+86

OpenStudy (amistre64):

that is our demand function;.....

OpenStudy (amistre64):

which can be rewritten as: D(x) = (2150-x)/25

OpenStudy (amistre64):

R * D(x) = total revenue and that has to be at least equal to 3750 to break even right?

OpenStudy (amistre64):

since R is the rent charged and x = rent in our demand function; lets do this.. x(2150-x) --------- >= 3750 25

OpenStudy (amistre64):

2150x - x^2 >= 93750 0 >= x^2 -2150x + 93750 +2150/2 is the x amount at the max :) 1075 = x

OpenStudy (amistre64):

lets see if that amount works out....

OpenStudy (amistre64):

2150-1075 ---------- = rooms rented 25 2150 1075 ----- 1075/25 = 43 43(1075)>= 3750 1045 43 ----- 4180 3135 ------- $44935 >= 3750...well, we make a profit at least :)

OpenStudy (amistre64):

44935 3750 ------ 41185 in profits

OpenStudy (amistre64):

we can trey that with 25 higher and 25 lower to see if we are right lol

OpenStudy (amistre64):

2150-1100 ---------- = rooms rented 25 2150 1100 ----- 1050/25 = 42 ...really? i had to figure it out that way lol...im an idiot :) of course 25 higher is one less lol

OpenStudy (amistre64):

1100 42 ----- 4400 2200 ------- 46200 ......................... see? 1050 44 ----- 4200 4200 ------ 46200

OpenStudy (amistre64):

DOH!!...now wheres my mistake lol

OpenStudy (amistre64):

1075 43 ----- 4300 3225 ------ 46225 ...thats better lol, i forgot how to multiply